OpenStudy (mendicant_bias):

(PDE)(Unknown) My instructor is posting (and testing us on) a lot of material outside of our textbook as standalone notes, and I don't understand what's going on in some of this, more info posted below.

OpenStudy (mendicant_bias):

The first thing is this right here: http://i.imgur.com/0eHCeCi.png I have no idea what I_n generally represents. Yeah, it's an integral, but what's the point of it? Does it have a name? Is this part of some general process? This is related to Gram-Schmidt, but my book once more doesn't cover it, and the sources I can find Gram-Schmidt mostly cover vector orthonormalization/orthogonalization, not *function* orthogonalization. @Kainui

OpenStudy (mendicant_bias):

The second thing is this, because all I have are these notes: How are f_1, f_2, f_3, and so on, related to f_n, between the previous and the following image? http://i.imgur.com/QiRiN8j.png

OpenStudy (mendicant_bias):

@iambatman , are you any good with this?

OpenStudy (mendicant_bias):

(And sorry, f_0, f_1, and f_2, not f_1, f_2, and f_3.)

OpenStudy (mendicant_bias):

@jim_thompson5910

OpenStudy (mendicant_bias):

Alright, I get what (in the second image) happened in I_0, now trying to figure out whatever is happening in I_n.

OpenStudy (dan815):

plug n=0

OpenStudy (mendicant_bias):

I'm understanding bits and pieces of what the expressi-yeah, I said I got that, n=0 isn't the issue, it's some of the other stuff. In I_n is where I'm a bit confused now, he has an expression, and he sets it equal to the integral of that expression's derivative plus something else, the original term evaluated from 0 to infinity.

OpenStudy (dan815):

where

OpenStudy (mendicant_bias):

Yeah, I just have....I_n. At I_n.

OpenStudy (dan815):

I_n=n*I_n-1

OpenStudy (dan815):

seen from integration by parts

OpenStudy (mendicant_bias):

I get integration by parts now on my own looking at that, but I don't understand what you just said, lmao. Thanks. I'm starting to get this.

OpenStudy (dan815):

\[I_n=n*I_{n-1}\]

OpenStudy (dan815):

so it cascades

OpenStudy (dan815):

\[I_n=n*I_{n-1}=n*(n-1)I_{n-2}=n*(n-1)(n-2)(n-3)...I_{n-n}\] and we know I_o =1

OpenStudy (dan815):

so n!

OpenStudy (mendicant_bias):

Yeah, I got it, it was the way you first wrote it without LaTeX; thanks so much, this makes sense now. I guess I'm still asking: How is this relevant to Gram-Schmidt, or is this? Let me just post the whole notes example to see if this makes sense in context, one sec.

OpenStudy (mendicant_bias):

(2nd and 3rd solution are omitted, just dealing with this right now) http://i.imgur.com/eBcqfEy.png

OpenStudy (dan815):

look at the inner product of the 2 functions and see if it is 0, that is the definition of functional orthogonality

OpenStudy (mendicant_bias):

Parts II and III seem very relevant, but I don't understand how part I is related to Gram-Scmhidt.

OpenStudy (mendicant_bias):

Yeah, just the first part. I mean. And there are three functions, does that mean (if they were orthogonal) I would be taking the inner product of f_0f_1, f_0f_2, and f_1f_2? It's still just the first part with the integrals, I don't see the relevance.

OpenStudy (mendicant_bias):

?

OpenStudy (mendicant_bias):

@dan815

OpenStudy (dan815):

yea

OpenStudy (mendicant_bias):

Yeah, just what I asked earlier, I fail to see the relevance of Part I to the Gram-Schmidt procedure. That's all I want to know atm, is how that is relevant. We found that I_n = n!, and how does that affect or change the process whatsoever.

OpenStudy (mendicant_bias):

(Or is it not apparently relevant, and the document was just...poorly named and labeled, overall, you think.)

OpenStudy (dan815):

it shows u relationship betwen f_ns

OpenStudy (dan815):

you just have an integral representation of n!

OpenStudy (dan815):

now u can see what n! means if n is not integer if such a thing exists

OpenStudy (mendicant_bias):

I've gotten this far, but I don't understand this last question. Either way, thank you very much, I'm going to have to think about this.

OpenStudy (dan815):

ok :)